OpenStudy (anonymous):

When f(x)=ln(2x^2+8)+0.1e^(2x^3), find f'(1)

OpenStudy (anonymous):

f'(x) = (4*x)/(2*x^2+8) + 0.6*x^2 e^(2*x^3) f'(1) = 4/10 + 0.6*e^2

OpenStudy (anonymous):

suppose u is a function (lnu)'=u'/u and (a^u)'=u'a^u dmhv solved from that way

OpenStudy (anonymous):

Thank you! :]